OpenStudy (anonymous):

Find The Value Of: 3 Times X Over The Square Root Of X Cubed.

OpenStudy (campbell_st):

\[\frac{3x}{\sqrt{x^3}}\] it can't be solved but it can be simplified \[\frac{3x}{\sqrt{x^2 \times x}} = \frac{3x}{x \sqrt{x}}\] cancel the common factor of x leaves \[\frac{3}{\sqrt{x}}\]

OpenStudy (anonymous):

THANK YOU!!(: